OpenStudy (anonymous):

through (5,4), parallel to y=4/5x+4

Mehek (mehek14):

parallel lines = same slope

OpenStudy (anonymous):

yes

Mehek (mehek14):

so that means the slope of the new line is 4/5 y = 4/5x + b we need to find b plug (5,4) into the equation

OpenStudy (anonymous):

would it be y=4/5x

Mehek (mehek14):

ye no y-ntercept

OpenStudy (anonymous):

i dont know why i put the x there i just noticed it sorry i didn't mean to Y= \[\frac{ 4 }{ 5 }\]

Mehek (mehek14):

\(y=\dfrac{4}{5}\)

Mehek (mehek14):

like that?

Mehek (mehek14):

it should be \(y=\dfrac{4}{5}x\)